∫_{a}^{b} f(x)dx = lim_{Δx→0} ∑_{i=1}^{n} f(x_{i})Δx

|x| = \{ \begin{matrix}x & if & x ≥ 0 \\ −x & if & x < 0 \end{matrix}

We say that the limit of f as x approaches a is L and write

lim_{x→a} f(x) = L

if, for every ϵ > 0, there exists a δ > 0 such that

|f(x) − L| < ϵ whenever |x − a| < δ

Define f : ℝ × ℝ → ℝ by

f(x, y) = \frac{1}{\sqrt{1 + x^{2} + y^{2}}} , x, y ∈ ℝ

The truth table for P ∨ Q is:

P Q P ∨ Q
T T T
T F T
F T T
F F F